# Controlling Counterfactual Harm in Decision Support Systems Based on Prediction Sets

### v1

@@ Appendix C Additional experimental results on counterfactual monotonicity
Appendix E Experiments under the interventional monotonicity assumption
In this section, we show the average accuracy $A(\lambda)$ achieved by a human participant using $\mathcal{C}_{\lambda}$, as predicted by the mixture of MNLs, against the average counterfactual harm upper-bound caused by $\mathcal{C}_{\lambda}$ for several $\alpha$ values on the strata of images with $\omega=80$ in Figure 7, with $\omega=95$ in Figure 8 and with $\omega=110$ in Figure 9. Here, each point corresponds to a different $\lambda$ value and its coloring indicates the empirical probability that $\lambda$ is included in the harm-controlling set $\Lambda^{\prime}(\alpha)$ given by Corollary 2, for a fixed $\alpha^{\prime}$ value across the random samplings of the test and calibration sets. For each classifier and strata of images, we select the $\alpha^{\prime}$ value maximizing the expected size of the harm-controlling set $\Lambda^{\prime}(\alpha)$. The results on all strata of images support the findings derived from the experiments assuming that the counterfactual monotonicity assumption holds. In addition, we observe that the gap between the lower-bound of the average counterfactual harm (that is equal to the average counterfactual harm under the counterfactual monotonicity assumption) and the upper-bound is often large.

### v2

@@ Appendix E Experimental results on counterfactual monotonicity with a more complex set-valued predictor
Appendix G Experiments under the interventional monotonicity assumption
In this section, we show the average accuracy $A(\lambda)$ achieved by a human expert using $\mathcal{C}_{\lambda}$, as predicted by the mixture of MNLs, against the average counterfactual harm upper-bound caused by $\mathcal{C}_{\lambda}$ for several $\alpha$ values on the strata of images with $\omega=80$ in Figure 13, with $\omega=95$ in Figure 14 and with $\omega=110$ in Figure 15. Here, each point corresponds to a different $\lambda$ value and its coloring indicates the empirical probability that $\lambda$ is included in the harm-controlling set $\Lambda^{\prime}(\alpha)$ given by Corollary 2, for a fixed $\alpha^{\prime}$ value across the random samplings of the test and calibration sets. For each classifier and stratum of images, we select the $\alpha^{\prime}$ value maximizing the expected size of the harm-controlling set $\Lambda^{\prime}(\alpha)$. The results on all strata of images support the findings derived from the experiments assuming that the counterfactual monotonicity assumption holds. In addition, we observe that the gap between the lower-bound of the average counterfactual harm (that is equal to the average counterfactual harm under the counterfactual monotonicity assumption) and the upper-bound is often large.